In the year 2006 it was discovered that global population was increasing at the annual rate of 1.7%. In what year will the global population be double?
Answer by scott8148(6628)

let x="year of doubling" ... 2=(1.017)^(x-2006) ... log2=(x-2006)(log1.017)

(log(2))/(log(1.017))=x-2006